For the following differential equations, find ag, 03, 04, 05, 06, and ay in terms of do and ay and write the answer y(x) = 4 just #3
Answer #1

nun a 6x6t 7 taxt... 다 ti let y = I a = a tantan²+ az x² + ax tapaht ay= a; + 2ax +39 0² +4 a gx² + 5agaat bagus + 7a 26 =y--- | 24 tоль 6 — 4 б + 4/%, то Ол = а. 5, +1 C = toy (2. бо 1 4 84 *0.5, Э % 4+ +19 s = 0 ta 2 84 0 : 2 с — 19 Х 12 * 1 1 J